### v2

# DeepMath-103K: A Large-Scale, Challenging, Decontaminated, and Verifiable Mathematical Dataset for Advancing Reasoning

Abstract
Reinforcement learning (RL) with large language models shows promise in complex reasoning. However, its progress is hindered by the lack of large-scale training data that is sufficiently challenging, contamination-free and verifiable. To this end, we introduce DeepMath-103K, a large-scale mathematical dataset designed with high difficulty (primarily levels 5-9), rigorous decontamination against numerous benchmarks, and verifiable answers for rule-based RL reward. It further includes three distinct R1 solutions adaptable for diverse training paradigms such as supervised fine-tuning (SFT). Spanning a wide range of mathematical topics, DeepMath-103K fosters the development of generalizable and advancing reasoning. Notably, models trained on DeepMath-103K achieve state-of-the-art results on challenging mathematical benchmarks and demonstrate generalization beyond math such as biology, physics and chemistry, underscoring its broad efficacy.
(a) Difficulty Levels of different datasets.
(b) Pass@1 Accuracies on AIME25.
Figure 1: (a) DeepMath-103K is challenging compared to existing datasets. (b) Results of DeepMath series models under zero RL and RL setting using DeepMath-103K.
1 Introduction
Reinforcement learning (RL) with large language models (LLMs) has demonstrated significant potential in complex mathematical reasoning (Guo et al., 2025; Hu et al., 2025; Zeng et al., 2025a; Liu et al., 2025). Despite this promise, the effective advancement of RL is constrained by existing training data. While numerous datasets are available, they fall short in several key aspects crucial for training advanced reasoning models: (1) insufficient difficulty (Figure 1a) to push the boundaries of current models (Dang & Ngo, 2025; Yu et al., 2025; Luo et al., 2025; Face, 2025; Hu et al., 2025), (2) contamination with standard benchmarks (appendix A), (3) a lack of verifiable answers essential for RL with verifiable rewards (RLVR) (Guo et al., 2025; Cobbe et al., 2021; Hendrycks et al., 2021b; Yu et al., 2024), or (4) an inadequate combination of these critical aspects at scale. Furthermore, many of existing datasets represent the recombination and filtration of common sources (such as AIME (MAA, a)) which contain already well-formatted data, thus lacking a substantial influx of novel and diverse problems from more varied but less structured sources (Dang & Ngo, 2025; Yu et al., 2025; Luo et al., 2025; Face, 2025; Hu et al., 2025).
To bridge this gap, we introduce DeepMath-103K, a large-scale mathematical dataset tailored for advancing reasoning via RLVR. DeepMath-103K distinguishes itself through several key features.
  • Challenging Problems: DeepMath-103K features a high concentration of challenging mathematical problems, with a difficulty distribution skewed towards higher levels ($\geq 5$) compared to existing open resources (Figure 1a).
  • Rigorous Decontamination: To ensure trustworthy evaluation, DeepMath-103K underwent a rigorous decontamination process against a comprehensive suite of benchmarks.
  • Verifiable Answers and Diverse Solutions: To enable rule-based reward functions in RLVR, every problem in DeepMath-103K includes a verifiable final answer that has been validated for easy extraction and verification via rules. Each problem is further enriched with three distinct R1 solutions (Guo et al., 2025), supporting diverse training paradigms such as SFT.
Beyond these core features, DeepMath-103K also differentiates itself in its raw data acquisition. The prevalent trend in existing open datasets often recombines readily available and well-formatted problems from common sources such as AIME (MAA, a). This approach does not create new problems, but re-collect existing ones, which leads to significant overlaps among different datasets. Recognizing the potential limitations and eventual exhaustion of common resources, DeepMath-103K draws its content from more diverse but less structured sources, notably including discussions from Math StackExchange [^1] https://math.stackexchange.com. The raw content from these sources is informal discourse and lacking a standard format. After a rigorous curation pipeline that transformed these discussions into a well-structured QA format, DeepMath-103K is characterized by its unique problem variety and diversity compared to existing datasets.
Consequently, models trained on DeepMath-103K achieve state-of-the-art (SOTA) results (Figure 1b):
  • Zero RL Training: Starting from the Qwen-2.5-(Math)-7B (Team, 2024), DeepMath-Zero-(Math)-7B shows pass@1 improvements of +12.7 (+23.0) on AIME24 and +12.1 (+19.1) on AIME25, establishing new SOTA performance.
  • RL Training: Initialized from instruction-tuned models, DeepMath variants also show substantial gains. DeepMath-1.5B, starting from R1-Distill-Qwen-1.5B (Guo et al., 2025), achieves pass@1 accuracy improvements of +7.9 on AIME24 and +6.0 AIME25. DeepMath-Omn-1.5B, built upon OpenMath-Nemotron-1.5B (Moshkov et al., 2025), reaches new SOTA pass@1 accuracies of 64.0 on AIME24 and 57.3 on AIME25, surpassing o1-mini (63.6 on AIME24) and low effort o3‑mini (60.0 on AIME24).
  • Generalizable Reasoning beyond Math: DeepMath series models also generalizes their reasoning abilities to broader domains, achieving best GPQA-Diamond (Rein et al., 2024) scores on biology, physics, and chemistry compared to the baselines.
These results underscore the value of DeepMath-103K as a resource for developing advanced reasoning models with broad applicability. The remainder of this paper is organized as follows:
  • § 2 presents an overview of DeepMath-103K, including its format, difficulty distribution, and topic covered;
  • § 3 details the data curation pipeline to construct DeepMath-103K, encompassing source analysis, decontamination, difficulty filtering, and robust answer verification
  • § 4 trains, evaluates and analyzes DeepMath series models that trained on DeepMath-103K.
To foster future research, we have released the DeepMath-103K dataset, along with the code and model weights, hoping to enable further exploration of advanced reasoning techniques and the development of robust and generalizable machine intelligence.
2 Overview of DeepMath-103K
Figure 2: A data sample from DeepMath-103K.
Each data sample in DeepMath-103K is intentionally structured to be comprehensive, supporting a variety of downstream applications in mathematical reasoning research. As illustrated in Figure 2, a single sample includes the following components:
  • Question: The mathematical problem statement.
  • Final Answer: A verifiable final answer, crucial for rule-based reward functions in RLVR.
  • Difficulty: A numerical difficulty score, which facilitates techniques like difficulty-aware training (e.g., curriculum learning) or adaptive compute allocation based on problem complexity (Wang et al., 2025b; Chen et al., 2024).
  • Topic: A hierarchical topic classification for the problem, enabling topic-specific analysis.
  • R1 Solutions: Three distinct reasoning paths generated by the DeepSeek-R1 model (Guo et al., 2025), suitable for diverse training paradigms such as SFT.
DeepMath-103K possesses several key characteristics that make it particularly suitable for advancing mathematical reasoning research:
Higher Difficulty
DeepMath-103K includes mathematical problems spanning difficulty levels 3 through 9. The core of the dataset consists of 95K challenging problems (levels 5-9) specifically curated for this research. To ensure broader difficulty coverage, this is augmented with an additional 8K problems (levels 3-5) sourced from SimpleRL (Zeng et al., 2025b). For comparison, we analyzed and labeled the difficulty levels of several existing datasets commonly used for RLVR training in math domain: Open-RS (Dang & Ngo, 2025), DAPO-17K (Yu et al., 2025), DSR-Preview (Luo et al., 2025), SITLL-3-RL (Chen et al., 2025), ORZ-129K (Hu et al., 2025), and Open-R1 (Face, 2025). Figure 1a illustrates the difficulty distributions across these datasets. As depicted, DeepMath-103K exhibits a significantly more challenging problem distribution, containing a substantially higher proportion of problems at difficulty levels 5 and above compared to the other benchmark datasets. This focus on higher difficulty is intended to push the reasoning limits of current models.
Rigorous Data Decontamination
Figure 3: Contamination rates of common mathematical and STEM benchmarks detected in the raw data sources before decontamination.
DeepMath-103K was constructed exclusively using the training splits of existing open resources, with careful avoidance of any known test set materials. However, our preliminary analysis revealed that these source data exhibits alarmingly high levels of contamination with problems from commonly used evaluation benchmarks. As illustrated in Figure 3, the contamination rates (defined as the percentage of benchmark test samples found within our raw data pool) are notably high: reaching 90% for AIME24 and AMC23, 76.6% for MATH500, 35.7% for Minerva Math, and 33.6% for OlympiadBench. Recognizing that these benchmarks are frequently employed for model evaluation, DeepMath-103K underwent a rigorous decontamination procedure. This process systematically identified and removed problems that overlap with these standard evaluation sets, ensuring the integrity and reliability of future benchmark results obtained using models trained on DeepMath-103K.
Broad Topical Diversity
Figure 4: Hierarchical breakdown of covered mathematical topics in DeepMath-103K.
Complementing its high difficulty and data integrity, a key characteristic of DeepMath-103K is its extensive topical diversity spanning the mathematical landscape. We categorized each problem using a hierarchical topic structure, following the methodology from Gao et al. (2024). As illustrated in Figure 4, this classification reveals that DeepMath-103K draws problems from a multitude of core mathematical areas. Its scope ranges from fundamental topics such as Prealgebra and Plane Geometry to sophisticated domains like Abstract Algebra (including Group Theory and Field Theory) and advanced Calculus (covering Differential Equations and Applications of Integrals, among others). This broad and deep topical foundation ensures that models trained on DeepMath-103K are exposed to a rich variety of mathematical concepts and problem-solving paradigms, thereby fostering the development of more robust and widely generalizable reasoning skills.
Data Novelty and Uniqueness
As mentioned in § 1, DeepMath-103K sources mostly from math forum, rather than common resources frequently adopted by other datasets. To evaluate the data novelty and uniqueness of DeepMath-103K, we performed the following analysis for all the datasets:
  1. 1.
    We first embedded all the samples using paraphrase-multilingual-MiniLM-L12-v2.
  2. 2.
    Samples with an embedding similarity greater than 0.98 were considered as the same samples.
Viewing each dataset as a set of embeddings, Figure 5 presents the number of unique elements in each set and the corresponding set sizes. DeepMath-103K contains 82.81K problems that are not found in others. This stark contrast highlights the data novelty and uniqueness of DeepMath-103K. We also plot their embedding distribution after t-SNE in Figure 6. ORZ-129K (Hu et al., 2025), Open-R1 (Face, 2025), SITLL-3-RL (Chen et al., 2025), DSR-Preview (Luo et al., 2025), and DAPO-17K (Yu et al., 2025), though curated independently, show very similar embedding distribution, while DeepMath-103K exhibits a distinctly different pattern. This observation supports our claim that existing datasets overlap with each other because of using common data sources and further demonstrate the data novelty and uniqueness of DeepMath-103K.
Figure 5: Unique and non-unique problem counts in DeepMath-103K compared to other datasets.
(a) DeepMath-103K
(b) ORZ-129K
(c) Open-R1
(d) STILL-3-RL
(e) DSR-Preview
(f) DAPO-17K
Figure 6: Embedding distributions of different datasets after t-SNE.
3 Construction of DeepMath-103K
Figure 7: The data curation pipeline for DeepMath-103K. Starting with an initial pool of 2,869K raw questions, successive stages of data decontamination, difficulty filtering (retaining levels $\geq$5), and answer verifiability filtering yield 95K problems. These are then combined with 8K problems from SimpleRL (Zeng et al., 2025b) to form the final DeepMath-103K dataset.
This section details the meticulous data curation process used to construct DeepMath-103K, illustrated in Figure 7. The process comprises four primary stages:
  1. 1.
    Source Analysis and Collection: Identifying and collecting mathematically challenging problems by analyzing the difficulty distributions of existing open data sources.
  2. 2.
    Data Decontamination: Rigorously decontaminating the collected data to remove potential overlaps with standard evaluation benchmarks, ensuring evaluation integrity.
  3. 3.
    Difficulty Filtering: Filtering the decontaminated problems based on difficulty, retaining only those assessed at level 5 or higher to focus on challenging content.
  4. 4.
    Answer Verification: Ensuring each curated problem possesses a verifiable final answer, consistently validated across multiple solution paths generated by DeepSeek-R1.
Overall, this curation pipeline ensures that DeepMath-103K is largely free from benchmark contamination and concentrates on challenging mathematical problems suitable for advanced reasoning model training. The entire procedure involved significant computational resources, requiring an expenditure of 138,000 US dollars in GPT-4o API fees and a total of 127,000 H20 GPU hours.
Figure 8: Difficulty distributions of various open mathematical datasets considered as potential sources.
Stage 1: Source Analysis and Collection.
To identify data sources rich in challenging problems, we began by analyzing the landscape of existing open mathematical reasoning datasets designed for SFT. These datasets utilize diverse collection methods. For instance, datasets such as MetaMathQA (Yu et al., 2024), dart-math-hard (Tong et al., 2024), and OpenMathInstruct-2 (Toshniwal et al., 2024a) primarily focus on augmenting problems and solutions derived from established datasets like GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b). In contrast, datasets like NuminaMath-CoT (LI et al., 2024), MMIQC (Liu et al., 2024), and WebInstructSub (Yue et al., 2024) source content more broadly from the web, gathering materials such as exercises and discussions from online platforms (e.g., Math Stack Exchange). We follow Gao et al. (2024) to estimate the difficulty distributions of these potential source datasets, as shown in Figure 8, which reveals distinct patterns: datasets derived from GSM8K and MATH (MetaMathQA, dart-math-hard, OpenMathInstruct-2), along with NuminaMath-CoT, exhibited distributions heavily skewed towards lower difficulty levels (levels 1-5). Conversely, datasets sourced more broadly from web content, specifically MMIQC and WebInstructSub, displayed significantly flatter distributions with a larger proportion of problems in the mid-to-high difficulty range (levels 5-9). Based on this finding, we selected Math StackExchange subsets from MMIQC and WebInstructSub as our primary data sources due to their higher concentration of challenging problems. We also included NuminaMath-CoT to enhance the topical diversity of the initial collection. After applying basic filtering, this selection process yielded a raw pool of 2,869K questions.
Stage 2: Data Decontamination.
As indicated by the high contamination rates observed in common benchmarks (Figure 3), a rigorous data decontamination process was crucial for ensuring the integrity of DeepMath-103K. We performed decontamination against a comprehensive suite of mathematics and STEM benchmarks, including MATH (Hendrycks et al., 2021b), AIME (MAA, a), AMC (MAA, b), Minerva Math (Lewkowycz et al., 2022), OlympiadBench (He et al., 2024), Omni-MATH (Gao et al., 2024), MathOdyssey (Fang et al., 2024), GAOKAO (Zhong et al., 2023), JEEBench (Arora et al., 2023), MMLU-STEM (Hendrycks et al., 2021a), CMATH (Wei et al., 2023), OlympicArena (Huang et al., 2024), GSM8K (Cobbe et al., 2021), and GPQA (Rein et al., 2024). We adopted the decontamination method proposed by Toshniwal et al. (2024a):
  1. 1.
    For each candidate question in our raw dataset, we employed embedding similarity search (using paraphrase-multilingual-MiniLM-L12-v2 (Reimers & Gurevych, 2019)) to identify the top-$k$ ($k=5$) most similar examples from the aggregated test sets of all targeted benchmarks.
  2. 2.
    Each candidate question was then compared against its top-$k$ retrieved benchmark examples using an LLM-Judge (Llama-3.3-70B-Instruct (Grattafiori et al., 2024)) to determine if they constituted identical questions or paraphrases. If any of these comparisons indicated a potential paraphrase or duplicate, the candidate question was discarded.
Table 1 illustrates the effectiveness of semantic decontamination compared to simple lexical matching. This approach aims to identify not only exact duplicates but also near-duplicates and paraphrased questions that might otherwise overlap with evaluation sets.
Table 1: Examples of contamination detected between the raw data pool and benchmarks using semantic comparison. Colors highlight conceptual or textual similarities.
Benchmark Raw Question Benchmark Question
AIME24 How many routes are there through from top left corner to bottom right in a 20x20 grid? I’m trying to solve this computer programming problem on Project Euler. I’ve seen a solution using nCr, where n = 40 and r = 20. Could someone explain to me how this work, please? Consider the paths of length 16 that follow the lines from the lower left corner to the upper right corner on an 8x8 grid. Find the number of such paths that change direction exactly four times, as in the examples shown below.
AMC23 Using only 3 paise, 5 paise, and 9 paise coins, what is the largest amount that cannot be paid in exact change? In the state of Coinland, coins have values 6,10, and 15 cents. Suppose x is the value in cents of the most expensive item in Coinland that cannot be purchased using these coins with exact change. What is the sum of the digits of x?
Stage 3: Difficulty Filtering.
Table 2: Examples of geometry problems retained by the difficulty filtering process (level $\geq$ 5).
Difficulty Problem
5 Four random points are placed in the plane, with each point’s horizontal and vertical coordinates uniformly distributed on the interval $(0,1)$. What is the expected largest size of a subset of these points that can form the vertices of a convex polygon?
6 A square has one side lying on the line $y=2x-17$ and two other vertices on the parabola $y=x^{2}$. Determine the minimum possible area of the square.”
7 Determine the sequence $s(k,n)$, which represents the number of sides of the intersection of a unit-radius regular polygon $P_{k}$ with $k$ sides and a rotating unit-radius regular polygon $P_{n}$ with $n\geq k$ sides, as the angle of rotation $\theta$ varies from 00 to $2\pi$. Provide the sequence $s(k,n)$ for all $n\geq k$.
8 Consider a convex n-gon $A_{1}A_{2}\cdots A_{n}$ inscribed in a unit circle. Determine the maximum value of the sum of the squares of all its sides and diagonals
9 Determine the maximal cardinality of a collection $\mathcal{C}$ of projective planes on $\omega$ such that no two distinct members of $\mathcal{C}$ are isomorphic. A set $L\subseteq\mathcal{P}(X)$ is a projective plane on $X\neq\emptyset$ if: 1. For any distinct $x,y\in X$, there is a unique $l\in L$ such that $x,y\in l$. 2. For any distinct $l,m\in L$, $|l\cap m|=1$. 3. There exist four distinct elements of $X$ such that no member of $L$ contains more than two of these four elements. Two projective planes $L$ and $M$ on $X$ are isomorphic if there is a bijection $\varphi:X\to X$ such that $l\in L$ if and only if $\varphi(l)\in M$.
Zeng et al. (2025a) highlights the importance of aligning RL training data difficulty with the target model’s reasoning capabilities, noting that powerful models benefit significantly from exposure to challenging problems. Building on this insight, our curation process for DeepMath-103K focuses on selecting problems that represent a significant reasoning challenge. To quantify difficulty, we adopted the approach detailed in Omni-MATH (Gao et al., 2024). We assigned a difficulty level to each decontaminated problem by prompting GPT-4o based on the annotation guidelines provided by the AoPS [^2] https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki:Competition_ratings. To ensure a robust estimate, we queried GPT-4o six times for each problem and averaged the resulting ratings to determine its final difficulty level. Subsequently, we applied a strict filtering criterion, retaining only those problems with an estimated difficulty level of 5 or higher. Table 2 showcases examples of geometry problems that passed this filtering stage, illustrating how increasing difficulty levels often correlate with greater conceptual depth and reasoning complexity.
Stage 4: Answer Verification.
The availability of verifiable final answers is crucial for enabling rule-based reward in RLVR, which helps mitigate reward hacking and has been instrumental in training successful reasoning models like DeepSeek-R1 (Guo et al., 2025). However, reliably constructing such answers presents two primary challenges:
  1. 1.
    Some open-ended questions inherently lack a easily verifiable final answer.
  2. 2.
    Certain answers are excessively complex (e.g., lengthy expressions or intricate notation), making them challenging or even infeasible for automated rule-based verification.
To address these issues, we implemented a rigorous two-stage verification process:
  1. 1.
    Question Filtering and Formatting: We used GPT-4o to process the raw questions. Problem types inherently unsuitable for verification were discarded. Questions phrased conversationally were rewritten into a standardized format seeking a single, specific numerical or symbolic answer.
  2. 2.
    Answer Verification via Consistency Check: For questions successfully passing the above step, we generated three distinct solution paths using DeepSeek-R1. A rule-based verifier then extracted the final answer from each of these generated solutions, as well as from the original source solution (when available). We enforced strict consistency: only problems where all extracted final answers were identical were retained in the final dataset.
This combined approach of question standardization and multi-solution answer consistency checking ensures that every problem included in DeepMath-103K possesses a final answer that is robustly verifiable using automated rules.
4 DeepMath Series Models
This section presents a comprehensive evaluation of the mathematical and general reasoning capabilities of our DeepMath series of models, which were trained on DeepMath-103K.
4.1 Setup
Training Paradigms
We employed two distinct RL training paradigms:
  • Zero RL: This paradigm involves training LLMs from their base (non-instruction-tuned) version using RL (Guo et al., 2025). We used group relative policy optimization (GRPO) (Shao et al., 2024) with fixes from Yu et al. (2025), and trained Qwen-2.5-(Math)-7B with a rule-based reward (+1 for correct final answer, -1 otherwise). Detailed settings are available in Footnote 3, and SFT results are presented in Appendix B.
  • RL: We also performed RL on instruction-tuned models that already possessing math reasoning ability. We explored this using R1-Distill-Qwen-1.5B (Guo et al., 2025) and OpenMath-Nemotron-1.5B (Moshkov et al., 2025).
Evaluation
Following Zeng et al. (2025a; b), we assessed the mathematical performance of our models on: MATH-500 (Hendrycks et al., 2021b), AMC 2023 (MAA, b), OlympiadBench (He et al., 2024), Minerva Math (Lewkowycz et al., 2022), AIME 2024-2025 (MAA, a), and the English subset of PolyMath (Wang et al., 2025a). To investigate the generalization of reasoning abilities beyond mathematics, we used the GPQA-Diamond benchmark, which covers biology, physics and chemistry (Rein et al., 2024). For all evaluations, we adopted pass@1 accuracy (averaged over 16 samples) as the metric, and fixed the decoding parameters to temperature=0.6, top_p=0.95, and max_tokens=32K. To ensure a fair comparison and eliminate variance caused by the evaluation script, we re-evaluated the performance of all baseline models under our evaluation settings.
4.2 Mathematical Reasoning Results
Table 3: Math reasoning performance. “DeepMath” denotes models trained on DeepMath-103K.
Model MATH AMC Olympiad Minerva AIME AIME Poly
500 23 Bench Math 24 25 Math
Proprietary Models
o1-mini 63.6
o3-mini (low effort) 60.0
Zero RL from Base Model
Qwen-2.5-7B (Team, 2024) 54.8 35.3 27.8 16.2 7.7 5.4 28.1
[…]
Qwen-2.5-Math-7B (Team, 2024) 46.9 31.9 15.8 15.5 11.2 4.4 22.7
[…]
RL from Instruct Models
R1-Distill-Qwen-1.5B (Guo et al., 2025) 84.7 72.0 53.1 36.6 29.4 24.8 39.9
[…]
OpenMath-Nemotron-1.5B (Moshkov et al., 2025) 91.8 90.5 70.3 26.3 61.3 50.6 56.8
[…]
The results presented in Table 3 collectively demonstrate the effectiveness of DeepMath-103K as a valuable resource for advancing the state-of-the-art in mathematical reasoning:
Zero RL Training on Base Model
DeepMath-Zero-7B and DeepMath-Zero-Math-7B, trained from the base Qwen-2.5-7B and Qwen-2.5-Math-7B models, demonstrate significant performance gains and achieve new SOTA results on all evaluated benchmarks. These results highlight the effectiveness of DeepMath-103K in enabling the training of powerful mathematical reasoners from scratch.
RL Training on Instruction-tuned Models
Fine-tuning instruction-tuned models with RLVR on DeepMath-103K also yields notable performance enhancements. DeepMath-1.5B, initialized from R1-Distill-Qwen-1.5B, achieves strong performance, particularly on AMC23 (82.3%) and OlympiadBench (61.8%). Similarly, DeepMath-Omn-1.5B, starting from OpenMath-Nemotron-1.5B, attains new SOTA results among 1.5B-scale models on all evaluated benchmarks, and even surpasses o1-mini and o3-mini (low effort). The consistent improvements observed across different instruction-tuned baselines further validate the utility of DeepMath-103K in boosting strong models.
4.3 Generalizable Reasoning beyond Mathematics
Table 4 presents the reasoning performance of DeepMath models on the GPQA-Diamond (Rein et al., 2024), which covers biology, physics, and chemistry. DeepMath series models achieve superior performance compared to other baseline, demonstrating a remarkable capacity to generalize their reasoning abilities acquired from mathematics to broader domains. We attribute this generalization to the data diversity and rigorous curation of DeepMath-103K. By sourcing less structured but more diverse data like Math StackExchange, DeepMath-103K yields a dataset with unique and diverse problems. Furthermore, the rigorous curation pipeline ensures both the challenge and the integrity of the data. This exposure to a wider variety of problem-solving scenarios and reasoning styles likely equips our models with more robust and transferable reasoning skills.
Table 4: Performance on the GPQA-Diamond benchmark.
Model Biology Physics Chemistry Overall
Zero RL from Base Model
Qwen-2.5-7B 33.6 27.8 21.4 25.3
[…]
Qwen-2.5-Math-7B 32.2 26.0 21.1 24.3
[…]
RL from Instruct Models
R1-Distill-Qwen-1.5B 13.5 36.2 4.4 19.1
[…]
OpenMath-Nemotron-1.5B 12.8 23.5 18.9 20.3
[…]
4.4 Analysis of Zero RL Using DeepMath-103K
Figure 9 presents an analysis of the characteristics observed during the zero RL training of DeepMath-Zero-7B. Specifically, Figure 9a illustrates the trend of rollout response length throughout the training process, while Figure 9b tracks the emergence of four cognitive behaviors following the method outlined by Gandhi et al. (2025) and Zeng et al. (2025a). The increasing trends in both response length and the manifestation of cognitive behaviors suggest a reproduction of the “aha moment” phenomenon observed in DeepSeek-R1 (Guo et al., 2025), which leads to a long reasoning model. Furthermore, Figure 9c shows the average response lengths of different models on the evaluated benchmarks. The notably longer response lengths exhibited by DeepMath-Zero-7B indicate that DeepMath-103K serves as a valuable resource for research on long reasoning models, particularly concerning phenomena such as over- and under-thinking (Chen et al., 2024; Wang et al., 2025b).
(a) Response length (training)
(b) Change of cognitive behaviors.
(c) Response length (test)
Figure 9: (a) DeepMath-103K is challenging compared to existing datasets. (b) Results of zero RL and RL using DeepMath-103K on AIME25.
5 Related Work
Datasets for advancing mathematical reasoning of LLM falls into three main lines corresponding to the three stages of LLM post-training: continue pre-training (CPT), SFT and RL. CPT aims to inject fundamental mathematical knowledge into LLMs with representative works like OpenWebMath (Paster et al., 2023), MathPile (Wang et al., 2024), InfiMM-Web-Math (Han et al., 2024), FineMath (Allal et al., 2025), and MegaMath (Zhou et al., 2025). SFT has been a foundational approach, utilizing datasets like MATH (Hendrycks et al., 2021b) and GSM8K (Cobbe et al., 2021) which provide problems with step-by-step solutions to teach models reasoning patterns. Subsequent efforts have focused on creating larger, harder and more diverse SFT datasets, such as MetaMathQA (Yu et al., 2024), OpenMathInstruct (Toshniwal et al., 2024b; a), NuminaMath-CoT (LI et al., 2024), MMIQC (Liu et al., 2024), dart-math-hard (Tong et al., 2024), and OpenMathReasoning (Moshkov et al., 2025). Recent progress in RLVR catalyzes datasets with verifiable reward, such as Open-R1 (Face, 2025), ORZ-129K (Hu et al., 2025), DSR-Preview (Luo et al., 2025), Open-RS (Dang & Ngo, 2025), DAPO-17K (Yu et al., 2025), and BigMath (Albalak et al., 2025). DeepMath-103K distinguishes itself by a unique blend of high difficulty, rigorous decontamination, and verifiable answers.
6 Conclusion
In this work, we introduce DeepMath-103K, a large-scale mathematical dataset specifically designed to advance the reasoning capabilities of LLMs through RLVR. DeepMath-103K distinguishes itself through its high concentration of challenging problems, rigorous decontamination against a wide range of benchmarks, and the inclusion of verifiable final answers and multiple diverse solutions for each problem. Our data curation pipeline leverages the richness of less structured mathematical forums, resulting in a dataset with significant novelty and diversity compared to existing resources. Our experiments demonstrate the substantial impact of DeepMath-103K. Models trained on this dataset, the DeepMath series, achieve new SOTA results on many mathematical benchmarks and exhibit remarkable generalization to domains beyond mathematics. By releasing the DeepMath-103K dataset, along with our code and model weights, we aim to provide a robust platform for the community to further explore and push the boundaries of advanced reasoning.
Appendix A Contamination Analysis of Existing Datasets
Figure 10: Number of contaminated samples in various datasets when compared against the MATH500 benchmark.
We performed a contamination analysis of several existing datasets, including ORZ-129K (Hu et al., 2025), DSR-Preview (Luo et al., 2025), DAPO-17K (Yu et al., 2025), Open-RS (Bansal et al., 2025), Open-R1 (Face, 2025), and DeepMath-103K. Our analysis focused on detecting potential contamination from the MATH500 (Hendrycks et al., 2021b), a commonly used benchmark. We employed a string-based comparison method, specifically identifying cases where the normalized indel similarity between a problem in the analyzed dataset and a problem in MATH500 exceeded 90%. This approach is notably more lenient than the rigorous semantic decontamination procedure used in the construction of DeepMath-103K (§ 3). However, the numbers of contaminated samples shown in Figure 10 reveal that most of the analyzed datasets exhibit some degree of contamination, with the exception of DeepMath-103K.
Appendix B SFT Results
As mentioned in § 2, each problem in DeepMath-103K includes three distinct R1-generated solutions, facilitating SFT. We fine-tuned Qwen-2.5-7B using either the first R1 solution or all three solutions. Table 5 shows that SFT on DeepMath-103K with one solution each problem also significantly enhances base model performance across all benchmarks, with multiple solutions yielding further gains. However, SFT still lags behind RL.
Table 5: Math reasoning performance after fine-tuning Qwen-2.5-7B via SFT. We also add DeepMath-Zero-7B as an RL counterpart for reference.
Model MATH500 AMC23 Olympiad Minerva AIME24 AIME25
Bench Math
Qwen-2.5-7B 54.8 35.3 27.8 16.2 7.7 5.4
[…]
Appendix C Training Details
We use verl as the training framework [^3] https://github.com/volcengine/verl. Configurations for training DeepMath series models are listed in Table 6.
Table 6: Configurations for training DeepMath series models.
Config DeepMath-Zero-7B DeepMath-Zero-Math-7B DeepMath-1.5B DeepMath-Omn-1.5B
lr 1e-6 1e-6 1e-6 1e-6
kl_coef 0.0 0.0 1e-3 1e-3
max_prompt_length 2K 1K 2K 2K
max_response_length 10K 3K 24K 24K
train_batch_size 512 512 128 128
ppo_mini_batch_size 32 32 64 64
clip_ratio_low 0.20 0.20 0.20 0.20
clip_ratio_high 0.28 0.28 0.27 0.27
temperature 1.0 1.0 0.6 0.6
rollout.n 16 16 16 18
overlong_buffer.len 2K 512 4K 4K
total_training_steps 500 500 1800 700
Appendix D Licenses for Existing Assets
Table 7: Licenses for existing assets.
Type Asset License
Data MMIQC (Liu et al., 2024) Apache 2.0
WebInstSub (Yue et al., 2024) Apache 2.0
NuminaMath-CoT (LI et al., 2024) Apache 2.0
Code verl (Sheng et al., 2024) Apache 2.0
NeMo-Skills (Toshniwal et al., 2024b) Apache 2.0
Model Qwen-2.5-7B Team (2024) Apache 2.0
R1-Distill-Qwen-1.5B (Guo et al., 2025) MIT
OpenMath-Nemotron-1.5B (Moshkov et al., 2025) CC BY 4.0
Appendix E Limitations and Broader Impacts
While DeepMath-103K advances mathematical AI, its difficulty assessment relies on LLM evaluations, potentially introducing bias. Topical diversity may not be perfectly balanced, and the dataset’s creation was computationally intensive. Our manual analysis reveals judgment and multiple-choice questions whose answers might be matched successfully via random guess. However, DeepMath-103K’s public release can lower the barrier for RL reasoning research, accelerate progress on challenging problems, improve benchmark reliability, and foster more generalizable AI.